OpenStudy (anonymous):

Find dy/dx by implicit differentation: 5+3x=sin(xy^7)

OpenStudy (anonymous):

what have you got so far? do you know how to differentiate implicitly?

OpenStudy (anonymous):

I don't even know how to start.. My professor has a very strong accent and I was completely lost during the lecture today.

OpenStudy (anonymous):

what is written xy power????

OpenStudy (anonymous):

\[5+3x=\sin(xy^7)\]

OpenStudy (anonymous):

ok....just wait for answer

OpenStudy (anonymous):

I don't want the answer. I want to know how to do it.

OpenStudy (anonymous):

If you want some help actually understanding implicit differentiation and not just an answer then check these out: http://www.khanacademy.org/math/calculus/differential-calculus/v/implicit-differentiation http://tutorial.math.lamar.edu/Classes/CalcI/ImplicitDIff.aspx

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

i find Paul's Online Notes and Khan Academy to be the best at teaching things online :) goodluck :)

OpenStudy (anonymous):

5+3x=sin(xy7) taking derivative on both sides 3=cos(xy7)(x.d/dx(y7).d/dx(y))+y) 3=cos(xy7)(x7y6.dy/dx+y) 3-ycos(xy7)=cos(xy7)(x7y6.dy/dx) dy/dx=3-ycos(xy7)/cos(xy7)7xy6